Income statement and balance sheet data for Great Adventures, Inc., are provided belovw GREAT ADVENTURES, INC. Income Statement For the Year Ended December 31, 2020 Revenues: Service revenue (clinic, racing, TEAM) Sales revenue (MU watches) $529,000 104,000 Total revenues $633,000 Expenses: Cost of goods sold (MU watches) Operating expenses Depreciation expense Interest expense Income tax expense 63,000 303,576 43,000 29,024 52,800 Total expenses 491,400 Net income $141,600GREAT ADVENTURES, INC. Balance Sheets December 31, 2020 and 2019 Increase () or Decrease (D) 2020 2019 Assets Current assets: Cash 161190) 292190 34,500 15,950 11,950 131,000 28,000 13,300 10,300 6,500 ) Accounts recelvable Other current assets Land 2.650 1,650 ) Long term assets: 200,000 ) 0 1,200,000 ) 0 Buildings 58,000 (66,500) 58,000 (23,500) 43.000 Total assets $1746,090 $ 217100 Liabilities and Stockholders Equity Current liabilities: Accounts payable Interest payable Income tax payable $10,950 680 $8,300 680 34,500 2,650 0) 18,300 ) 392,440 ) Long-term liabilities: Notes payable 418,940 26.500 30,000 117120 Stocknolders equity Common stock Paid-n capital Retained earnings Treasury stock 100,000 1) 1106,0000) 17,600 ) 0 108,000) () 1106,000 134720 Total liabilities and stockholders equity $1,746,090 $ 217100 As you can tell from the financial statements, 2020 was an especially busy year. Tony and Suzie were able to use the $1.2 million received from the issuance of 100,000 shares of stock to hire a construction company for $1 million to build the cabins, dining facilities, ropes course, and the outdoor swimming pool. They even put in a baby pool to celebrate the birth of their firstborn son, little Venture Matheson. Assume all sales and services are on credit.2. Calculate the following profitability ratios for 2020. (Round your answers to 2 decimal places.) a. Gross profit ratio (on the MU watches) b. Return on assets c. Pofit margin d. Asset turnover e. Return on equity times
Answer #1

Answer 1(a) - Calculate Receivables Turnover Ratio -

Receivables Turnover Ratio = Net Credit Sales/Average Net Accounts Receivable

= 633,000/(34,500+28,000/2)

= 633,000/31,250

= 20.25 times

Answer 1(b) - Calculate Average Collection period -

Average collection period = 365/Receivable turnover ratio

=365/20.25 = 18.02 days

Answer 1(c) - Calculate Inventory turnover ratio -

Inventory turnover ratio = Cost of Goods Sold/Average Stock

= 63,000/ (15,950+ 13,300/2)

= 63,000/14,625

= 4.30 times

Answer 1(d) - Calculate average days in inventory -

average days in inventory = 365/ Average inventory ratio

=365/4.30

= 84.88 days

Answer 1(e) - Calculate Current ratio = Current assets/Current Liabilities

=Current assets - 292,190+34,500+15,950+11,950 = 354,590

=Current Liabiliites - 10,950+680+52,800 = 64,430

Current Ratio = 354,590/64,430 = 5.50 times

Answer 1(f) - Acid test Ratio = Current Asset- inventory/Current Liabilities

= 354,590 - 15,950/64,430

= 338,640/64,430

=5.25 times

Answer 1(g)- Calculate debt to equity ratio = Debt/total equity

=Debt - 64,430+418,940= 483,370

=Equity - 130,000+1,106,000+134,720+(108,000) = 12,62,720

=483,370/12,62,720

=38.28%

Answer 1(h) - Calculate Times interest earned ratio = EBIT/ Interest expense

EBIT = 141,600 +29,024+52,800 = 223,424

=223424/29,024

= 7.69 times

Answer 2(a) - Calculate Gross Profit Ratio = Gross Profit/Net Sales X 100

= Gross Profit = Net Sales (of MU watches only) - COGS

= 104,000-63,000 = 41,000

= 41,000/104,000

=39.42%

Answer 2(b) - Calculate Return on Assets = Net income/ Average Total Assets X 100

=141,600/(1,746,090/2) X 100

= 141,600/873,045 X 100

= 16.21%

Answer 2(c) - Calcuate Profit Margin = Net income/ Net sales X 100

=141,600/633,000

= 22.36%

Answer 2(d) - Calculate Asset Turnover - Net Sales/Average Total assets

=633,000/(1,746,090/2)

= 0.72 times

Answer 2(e) - Calculate Return on Equity = Net income/ Shareholder's equity

=141,600/ (130,000+1,106,000+134,720+(108,000)

= 141,600/1,262,720

=11.21%